# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 4. Preliminaries on Random Walks
Recall that $S$ be a finitely symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense, and $\chi_{S}$ be the uniform probability measure supported on $S$. The following lemmas are quantitative statements about non-concentration of self-convolutions of $\chi_{S}$ in proper subvarieties.
For the discussion in the following in this section, we prefix a primitive linear form
$L\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right),\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)=X_{1}x_{1}+Y_{1}y_{1}+Z_{1}z_{1}+W_{1}w_{1}+X_{2}x_{2}+Y_{2}y_{2}+Z_{2}z_{2}+W_{2}w_{2},$
i.e., $\text{gcd}(X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2})=1$.
@@
Let $S$ be a finite symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense in $\text{SL}_{2}\times\text{SL}_{2}$. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c}.$
@@
There are constants $10<c_{S},c<$ 1 depending only on $S$ such that the following holds. Let $Q\in\mathbb{Z}_{+}$ large enough and $\xi=(\xi_{1},\xi_{2}),\eta=(\eta_{1},\eta_{2})\in\text{Mat}_{2}(\mathbb{Z})\times\text{Mat}_{2}(\mathbb{Z})$ satisfy
$\text{Tr}(\xi_{1})=\text{Tr}(\xi_{2})=\text{Tr}(\eta_{1})=\text{Tr}(\eta_{2})=0$
and either for every $p|Q$, $\pi_{p}(\xi_{1}),\pi_{p}(\xi_{2})\neq 0$, or for every $p|Q$, $\pi_{p}(\eta_{1}),\pi_{p}(\eta_{2})\neq 0$. Then for $l>c_{S}\log Q$ and any $n\in\mathbb{Z}$,
@@
In the $\mathrm{ASL}_{2}$ case, we also have the following analogue about non-concentration in proper subvarieties.
@@
Let $S$ be a finite symmetric set on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense in $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
@@
We focus on proving Proposition 4.1; Proposition 4.2 and Proposition 4.4 follow in a similar way (see also Proposition 4.1 in [BG09]). We first show that Proposition 4.1 follow from
@@
There are constants $10<c_{S},c_{0}<$ 1 depending only on $S$ such that for any $Q\in\mathbb{Z}_{+},n\in\mathbb{Z}$ and for $l\in\mathbb{Z}_{+}$ with $1\ll_{S}l<c_{S}\log Q$, we have
$\chi_{S}^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L(g)\equiv n\pmod{{Q}}\})<e^{-c_{0}l}.$ (4.6)
@@
Let $c_{S},c_{0}$ be the constants given by Lemma 4.5.
Let $l_{0}=[c_{S}\log Q]$ and write $\chi_{S}^{(l)}=\chi_{S}^{l_{0}}*\chi_{S}^{(l-l_{0})}$. For any $g^{\prime}$ in the support of $\chi_{S}^{l-l_{0}}$, let $L_{g^{\prime}}(g)=L(gg^{\prime})$. Clearly, $L_{g^{\prime}}$ is also primitive, so Lemma 4.5 is applicable to $L_{g^{\prime}}$. Therefore,
@@
$=$ $\sum_{g^{\prime}\in\Lambda}\chi_{S}^{(l_{0})}\left(\left\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L_{(g^{\prime})^{-1}}(g)\equiv n(\operatorname{mod}Q)\right\}\right)\chi_{S}^{(l-l_{0})}(g^{\prime})$
@@
Proposition 4.1 is thus proved with $c=c_{0}c_{S}$
Now we focus on proving Lemma 4.5.
@@
By Theorem A , there exists an absolute constant $10<\lambda<$ 1 which is the upper bound for all eigenvalues of the family of the operators $T_{p}:l_{0}^{2}(\Lambda)\rightarrow l_{0}^{2}(\Lambda)$ defined at (2.1). So
$\|\chi_{S}^{(l)}-\dfrac{1}{|\pi_{p}(\Gamma)|}\boldsymbol{1}_{\pi_{p}(\Gamma)}\|_{2}\leq\lambda^{l}$ (4.8)
for all prime $p$. Then since $|\pi_{p}(\Gamma)|\approx p^{6}$, if $l>\frac{6\log p}{(\log 1/\lambda)}$,
@@
by counting points in the subvariety of codimension $1$ determined by a primitive linear equation. So we are done by picking any prime $p\in[e^{\frac{l}{8}\log\frac{1}{\lambda}},e^{\frac{l}{7}\log\frac{1}{\lambda}}]$, which exists by the Prime Number Theorem when $l$ is sufficiently large. This finishes the proof of Lemma 4.7. ∎
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ without common zeros in $\mathbb{C}^{n}$ with $n\geq 2,\operatorname{deg}\mathcal{P}_{j}\leqslant d,d\geqslant 3$, $h\left(\mathcal{P}_{j}\right)\leqslant h$. Here $h(\mathcal{P})=\log\max\limits_{i}\left|a_{i}\right|$ for $\mathcal{P}(x)=a_{0}+a_{1}x+a_{2}x^{2}+\cdots+a_{n}x^{n}$ is the logarithm height function of polynomials. Then there is an integer $D\in\mathbb{Z}_{+}$, polynomials $\mathcal{Q}_{1},\ldots,\mathcal{Q}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ such that
@@
Write $Q=\prod_{i\in I}p_{i}^{n_{i}}$. We divide our proof into two cases $n=0$ and $n\neq 0$. We first deal with the case $n=0$.
By the primitivity of $L$, for each $p|Q$, at least one of
$t\in\mathcal{A}:=\{X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2}\}$
@@
$Q_{t}=\prod_{\begin{subarray}{c}p_{i}^{n_{i}}||Q\\ \text{gcd}(p_{i},t)=1\end{subarray}}p_{i}^{n_{i}}$
@@
Assume $t=X_{1}$ without loss of generality. Then $Q^{\prime}||Q$ and $(Q^{\prime},X_{1})=1$.
Now let
$\|(g_{1},g_{2})\|=\max\{\text{absolute values of coefficients of }g_{1}\text{ and }g_{2}\}.$
Let $C_{1}$ be an upper bound of $\|g\|$ for all $g\in\operatorname{supp}\chi_{S}$. Define
@@
$\chi_{S}^{(l)}(\mathcal{G})<e^{-cl}.$
for some $c>0$. For each $\gamma=\left(\begin{pmatrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{pmatrix},\begin{pmatrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{pmatrix}\right)\in\mathcal{G}$ we introduce a linear polynomial $f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Q}[\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}]$ as follows:
$f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+{\tilde{Y}}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}$ (4.10)
@@
$X_{1}f_{\gamma}\left(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}}\right)\equiv 0\ (\operatorname{mod}{Q^{\prime}})\text{ for all }\gamma\in\mathcal{G}.$ (4.11)
Here $\bar{X_{1}}$ is the multiplicative inverse of $X_{1}$ mod $Q^{\prime}$.
Also, by definition of $\mathcal{G}$, the coefficients of $f_{\gamma}$, namely entries of $\gamma$, are bounded by $(2C_{1})^{l}$. Hence,
$h\left(f_{g}\right)<2l\log C_{1},$
We claim that there is a common zero $(\tilde{Y},\tilde{Z},\tilde{W},\tilde{A},\tilde{B},\tilde{C},\tilde{D})\in\mathbb{C}^{7}$ to the following system of equations:
$f_{\gamma}(Y,Z,W,A,B,C,D)=0\quad\text{ for all }\gamma\in\mathcal{G},$ (4.12)
@@
Assume the claim fails to hold. We invoke Theorem C with $n=7,d=3,h=2l\log C_{1}$. It follows that there is an integer $M\in\mathbb{Z}_{+}$ and polynomials $\varphi_{l}$ of degree at most $b=7\times 15\times 3^{7}$ satisfying
$M=\sum_{l=1}^{N}F_{l}\varphi_{l}$ (4.13)

### v4

@@ 4. Preliminaries on Random Walks
Recall that $S$ is a finite symmetric set in ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense, and $\chi_{S}$ be the uniform probability measure supported on $S$. In this section we record some quantitative statements on non-concentration of iterated convolutions of $\chi_{S}$ in certain subvarieties. Proposition 4.1 and Proposition 4.3 concern linear subvarieties, and Proposition 4.2 is on certain quadratic subvarieties.
@@
Let $S$ be a finite symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense in $\text{SL}_{2}\times\text{SL}_{2}$. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $0<c_{S},c<1$ depending only on $S$ such that for any sufficiently large $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$, and any $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c},$
where $L$ is a general primitive linear form on $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$, i.e.,
$L\left(\left(\begin{matrix}x_{0}&x_{1}\\ x_{2}&x_{3}\end{matrix}\right),\left(\begin{matrix}x_{4}&x_{5}\\ x_{6}&x_{7}\end{matrix}\right)\right)=X_{0}x_{0}+X_{1}x_{1}+X_{2}x_{2}+X_{3}x_{3}+X_{4}x_{4}+X_{5}x_{5}+X_{6}x_{6}+X_{7}x_{7},$
with $\text{gcd}(X_{0},X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},X_{7})=1$.
@@
Let $S$ be given as in Proposition 4.1. There are constants $0<c_{S},c<1$ depending only on $S$ such that the following holds. Suppose $\xi=(\xi_{1},\xi_{2}),\eta=(\eta_{1},\eta_{2})\in\text{Mat}_{2}(\mathbb{Z})\times\text{Mat}_{2}(\mathbb{Z})$ satisfy
$\text{Tr}(\xi_{1})=\text{Tr}(\xi_{2})=\text{Tr}(\eta_{1})=\text{Tr}(\eta_{2})=0,$
Then for any sufficiently large $Q\in\mathbb{Z}_{+}$ with either $\text{gcd}(Q,\text{entries of }\xi_{1}\text{ and }\eta_{1})=1$ or $\text{gcd}(Q,\text{entries of }\xi_{2}\text{ and }\eta_{2})=1$, for any $l>c_{S}\log Q$ and for any $n\in\mathbb{Z}$,
@@
In the $\mathrm{ASL}_{2}$ case, we need
@@
Let $S$ be a finite symmetric set on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $0<c_{S},c<1$ depending only on $S$ such that for any sufficiently large $Q\in\mathbb{Z}_{+}$, any $l>{c_{S}}{\log Q}$ and any $n\in\mathbb{Z}$, we have
@@
Remark 4.4.
In Proposition 4.1 and Proposition 4.3, it is important that the implied constant $c$ is independent of $L$ and $n$.
We focus on proving Proposition 4.1. Proposition 4.2 and Proposition 4.3 follow in a similar way. Our method is similar to the proof of Proposition 4.1 in [BG09].
We first show that Proposition 4.1 follows from
@@
There are constants $0<c_{S},c_{0}<1$ depending only on $S$ such that for any $Q\in\mathbb{Z}_{+},n\in\mathbb{Z}$ and for any $1\ll_{S}l\leq c_{S}\log Q$, we have
$\chi_{S}^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\})<e^{-c_{0}l}.$ (4.6)
@@
Let $c_{S},c_{0}$ be the constants given by Lemma 4.5. Let $l_{0}=[c_{S}\log Q]$ and write $\chi_{S}^{(l)}=\chi_{S}^{l_{0}}*\chi_{S}^{(l-l_{0})}$. For any $g^{\prime}$ in the support of $\chi_{S}^{l-l_{0}}$, let $L_{g^{\prime}}(g)=L(gg^{\prime})$. Clearly, $L_{g^{\prime}}$ is also primitive, so Lemma 4.5 is applicable to $L_{g^{\prime}}$. Therefore,
@@
$=$ $\sum_{g^{\prime}\in\Lambda}\chi_{S}^{(l_{0})}\left(\left\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L_{g^{\prime}}(g)\equiv n(\operatorname{mod}Q)\right\}\right)\chi_{S}^{(l-l_{0})}(g^{\prime})$
@@
Proposition 4.1 is thus proved with $c=c_{0}c_{S}$. ∎
Now we focus on proving Lemma 4.5. We begin with
@@
By Theorem A, there exists an absolute constant $0<\lambda<1$ which is the upper bound for all eigenvalues of the family of the operators $T_{p}:l_{0}^{2}(\Lambda)\rightarrow l_{0}^{2}(\Lambda)$ defined at (2.1). So
$\|\chi_{S}^{(l)}-\dfrac{1}{|\pi_{p}(\Gamma)|}\boldsymbol{1}_{\pi_{p}(\Gamma)}\|_{2}\leq\lambda^{l}\|\chi_{S}\|_{2}={\lambda^{l}}{|S|^{-\frac{1}{2}}}$ (4.8)
for all prime $p$. Since $|\pi_{p}(\Gamma)|\sim p^{6}$, if $l>\frac{6\log p}{\log(1/\lambda)}$,
@@
by counting points in subvarieties of codimension $1$ determined by primitive linear equations. So we are done by picking any prime $p\in[e^{\frac{l}{8}\log\frac{1}{\lambda}},e^{\frac{l}{7}\log\frac{1}{\lambda}}]$, which exists by the Prime Number Theorem when $l$ is sufficiently large. This finishes the proof of Lemma 4.7. ∎
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ without common zeros in $\mathbb{C}^{n}$ with $n\geq 2,\operatorname{deg}\mathcal{P}_{j}\leqslant d,d\geqslant 3$, $h\left(\mathcal{P}_{j}\right)\leqslant h$. Here $h(\mathcal{P})$ is the logarithmic of the modulus of the coefficient of $\mathcal{P}$ of largest absolute value. Then there is an integer $D\in\mathbb{Z}_{+}$, polynomials $\mathcal{Q}_{1},\ldots,\mathcal{Q}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ such that
@@
Write $Q=\prod_{i\in I}p_{i}^{n_{i}}$. We divide our proof into two cases $n=0$ and $n\neq 0$.
Case 1: $n=0$. By the primitivity of $L$, for each $p|Q$, at least one element in
$\mathcal{A}:=\{X_{0},X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},X_{7}\}$
@@
$Q_{t}=\prod_{\begin{subarray}{c}p_{i}^{n_{i}}\|Q\\ \text{gcd}(p_{i},t)=1\end{subarray}}p_{i}^{n_{i}}.$
@@
Assume $t=X_{0}$ without loss of generality. Then $Q^{\prime}\|Q$ and $(Q^{\prime},X_{0})=1$.
[…]
$\|(g_{1},g_{2})\|=\max\{\|g_{1}\|,\|g_{2}\|\}.$
Let $N_{1}$ be an upper bound of $\|g\|$ for all $g\in\operatorname{supp}\chi_{S}$.
Define
@@
$\chi_{S}^{(l)}(\mathcal{G})<e^{-cl},$ (4.10)
for some $c>0$.
For each $\gamma=\left(\begin{pmatrix}\gamma_{0}&\gamma_{1}\\ \gamma_{2}&\gamma_{3}\end{pmatrix},\begin{pmatrix}\gamma_{4}&\gamma_{5}\\ \gamma_{6}&\gamma_{7}\end{pmatrix}\right)\in\mathcal{G}$, we introduce a linear polynomial
$f_{\gamma}(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7})\in\mathbb{Q}[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7}]$
as follows:
$f_{\gamma}(x_{1},\cdots,x_{7})=\gamma_{0}+\sum_{i=1}^{7}\gamma_{i}x_{i}$ (4.11)
@@
$f_{\gamma}\left(X_{1}\overline{X_{0}},X_{2}\overline{X_{0}},X_{3}\overline{X_{0}},X_{4}\overline{X_{0}},X_{5}\overline{X_{0}},X_{6}\overline{X_{0}},X_{7}\overline{X_{0}}\right)\equiv 0\ (\operatorname{mod}{Q^{\prime}})$ (4.12)
for all $\gamma\in\mathcal{G}.$ Here $\overline{X_{0}}$ is the multiplicative inverse of $X_{0}$ mod $Q^{\prime}$.
Also, by the definition of $\mathcal{G}$, the coefficients of $f_{\gamma}$, namely the entries of $\gamma$, are bounded by $(2N_{1})^{l}$. Hence,
$h\left(f_{g}\right)<2l\log N_{1},$
Claim: There is a common zero $(\tilde{x}_{1},\cdots,\tilde{x}_{7})\in\mathbb{C}^{7}-\{\vec{\bf 0}\}$ to the following system of equations:
$f_{\gamma}({x}_{1},\cdots,{x}_{7})=0,\quad\gamma\in\mathcal{G}.$ (4.13)
Hence $\mathcal{G}$ is contained in some proper subvariety of $\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})$. Note that the size of $\mathcal{G}$ is bounded by $|S|^{l}$.
Assume the claim fails to hold. We invoke Theorem C with $n=7,d=3,h=2l\log N_{1}$ and $N\leq|S|^{l}$. It follows that there is an integer $M\in\mathbb{Z}_{+}$ and polynomials $\varphi_{\gamma}$ of degree at most $b=7\times 15\times 3^{7}$ satisfying
$\sum_{\gamma\in\mathcal{G}}f_{\gamma}\varphi_{\gamma}=M$ (4.14)
with
$0<\log M,h\left(\varphi_{\gamma}\right)<\mathfrak{X}(7)3^{59}((\log N_{1}+\log(2|S|))l+3\log 3)<N_{2}l$ (4.15)
where $N_{2}=3^{61}\mathfrak{X}(7)\log(2N_{1}|S|)$.
Now we take $(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6},y_{7})\in\mathbb{Z}^{7}$ such that
$(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6},y_{7})\equiv(X_{1}\overline{X_{0}},X_{2}\overline{X_{0}},X_{3}\overline{X_{0}},X_{4}\overline{X_{0}},X_{5}\overline{X_{0}},X_{6}\overline{X_{0}},X_{7}\overline{X_{0}})(\operatorname{mod}Q^{\prime}).$
It follows that
$\sum_{\gamma\in\mathcal{G}}f_{\gamma}(y_{1},\cdots,y_{7})\varphi_{\gamma}(y_{1},\cdots,y_{7})-M\equiv 0(\operatorname{mod}Q^{\prime}).$
@@
Since the linear system (4.13) admits a solution and the coefficients of $f_{\gamma}$ are all integral, it must admit a rational solution $(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3},\tilde{x}_{4},\tilde{x}_{5},\tilde{x}_{6},\tilde{x}_{7})$. In other words,
$\gamma_{0}+\sum_{i=1}^{7}\gamma_{i}\tilde{x}_{i}=0,\forall\gamma\in\mathcal{G}$ (4.16)
If at least one element $\gamma\in\mathcal{G}$ has $\gamma_{0}\neq 0$, then $(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3},\tilde{x}_{4},\tilde{x}_{5},\tilde{x}_{6},\tilde{x}_{7})\neq\vec{\bf{0}}$. Rescaling to get rid of the common denominator in (4.16), we obtain $(v_{0},v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7})\in\mathbb{Z}^{8}-\vec{\bf 0}$ such that $\text{gcd}(v_{0},v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7})=1$, and for all $\gamma\in\mathcal{G}$, we have
$\sum_{i=0}^{7}\gamma_{i}v_{i}=0.$
Apply Lemma 4.7 to the linear form $\tilde{L}$ determined by the constants $v_{1},v_{2},v_{0},v_{3},v_{4},v_{5},v_{6},v_{7}$: For $1\ll_{S}l<c_{S}\log Q$,
$\chi_{S}^{(l)}(\mathcal{G})\leq\$$\chi_{S}^{(l)}\left(\left\{\gamma\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})\mid\tilde{L}(\gamma)=0\right\}\right)$
[…]
where $c_{0}$ is the constant given in Lemma 4.7.
If all $\gamma\in\mathcal{G}$ have $\gamma_{0}=0$, then $\mathcal{G}$ lies in a linear subvariety, and (5.30) is given by Lemma 4.7. We have thus fully proved the case of Lemma 4.5 when $n=0$.
Case 2: $n\neq 0$. Write $Q_{1}=(Q,n)$, $n=Q_{1}n_{1}$ and $Q=Q_{1}Q_{2}$. We clearly have $(n_{1},Q_{2})=1$. We analyse the cases $Q_{1}$ small and $Q_{1}$ large separately.
Case 2.1: $Q_{1}<Q_{2}^{\frac{1}{3N_{3}}}$, where the constant $C_{3}$ is given at (4.19). Consider the following set:
$\mathcal{G}^{\prime}=\left\{g\in\text{supp }[\chi_{S}^{(l)}]\mid L(g)-Q_{1}n_{1}\equiv 0\ (\operatorname{mod}{Q_{2}})\right\}.$
For each $\gamma=\left(\left(\begin{matrix}\gamma_{0}&\gamma_{1}\\ \gamma_{2}&\gamma_{3}\end{matrix}\right),\left(\begin{matrix}\gamma_{4}&\gamma_{5}\\ \gamma_{6}&\gamma_{7}\end{matrix}\right)\right)\in\mathcal{G}^{\prime}$, define
$f_{\gamma}(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7})=\sum_{i=0}^{7}\gamma_{i}x_{i}-Q_{1}.$
If $\{f_{\gamma},\gamma\in\mathcal{G}\}$ has no common zero, then there exist integral polynomials $\varphi_{\gamma}$ and $M\in\mathbb{Z}_{+}$ such that
$\sum_{\gamma\in\Gamma}f_{\gamma}\varphi_{\gamma}=M.$ (4.17)